How many liters of water should be added to 20 liters of a 30% sulfuric acid solution to reduce it to a 15% solution?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How many liters of water should be added to 20 liters of a 30% sulfuric acid solution to reduce it to a 15% solution?
==============
Add w liters
20*30 + w*0 = (20+w)*15

Answer by stanbon(75887)

How many liters of water should be added to 20 liters of a 30% sulfuric acid solution to reduce it to a 15% solution?
-----
water + water = water
0.7*20 + 1*x = 0.85(20+x)
------
70*20 + 100x = 85*20 + 85x
-------
15x = 15*20
x = 20 liters (amt. of water to add)
-------------
Please let me know if this helps.
-------------
Cheers,
Stan H.
---------------